Using historical records, a manufacturing firm has developed the following probability distribution for the number of days required to get components from its suppliers. The distribution is shown here, where the random variable x is the number of days.

x P(x)
4 0.17
5 0.38
6 0.31
7 0.105
8 0.035

a. The average lead time is ______ days. (Type the entire number, do not round.)

b. The coefficient of variation is ______ %. (Round to two decimal places as needed. Do not round. Do not include % in your answer.)
Answer #1
x P(X=x) xP(x) x2P(x)
4 0.17000 0.68000 2.72000
5 0.38000 1.90000 9.50000
6 0.31000 1.86000 11.16000
7 0.10500 0.73500 5.14500
8 0.03500 0.28000 2.24000
total 5.4550 30.7650
E(x) =μ= ΣxP(x) = 5.4550
E(x2) = Σx2P(x) = 30.7650
Var(x)=σ2 = E(x2)-(E(x))2= 1.0080
std deviation=         σ= √σ2 = 1.0040
Coefficient of variation=(σ/µ)*100 = 18.40

a)

average lead time is =5.455

b) coefficient of variation is =18.40 %